OpenStudy (anonymous):

Eight coins are placed in a line on a level ground. The distance b/n any two consecutive coins is 10m. How far must a person travel to bring them one by one to a basket placed 10m behind the first coin. Assume that the basket is on the line of the coins and the person walks from the basket.

OpenStudy (anonymous):

thus for the first coin (as he starts from the basket each time )he travels 2*10 and for the second coin 2*2*10 and so on .. thus the required distance covered is 2*10*(1+2+3+...8)=2*10*8*9/2=720 m